OpenStudy (anonymous):

9= 12+ 5x - (6x+7) @jackellyn can you help me some more?

OpenStudy (mathstudent55):

First, get rid of the parentheses. To do that, distribute the negative sign to the left of the parentheses.

OpenStudy (jackellyn):

So distribute the -1 with the parenthesis first, then combine like terms and solve like we diid before

OpenStudy (anonymous):

I know we did this before. I guess it just helps to go over it a few times for me. im sorry if im a nuisance.

OpenStudy (jackellyn):

I am here to help! :) Show me the first step of distributing

OpenStudy (anonymous):

12+ 5x- (-6) + (-7) is that right?

OpenStudy (jackellyn):

don't forget the 9, your distribution is correct but the -6 has an x

OpenStudy (anonymous):

oh ya. forgot about the 9 haha. ok so next its 9+12+ (-7)= 5x- (-6x) is that right?

OpenStudy (mathstudent55):

No. 9 = 12+ 5x - (6x+7) A negative sign to the left of parentheses is the same as -1 multiplying the parentheses. 9 = 12 + 5x -1(6x + 7) Now you distribute the -1 throughout the terms inside the parentheses. 9 = 12 + 5x -1(6x) + (-1)(7) 9 = 12 + 5x - 6x - 7 Now collect like terms on the right side.

OpenStudy (jackellyn):

@samissmurfette , @mathstudent55 is right. Do you see all the steps he did? Now combine like terms, be careful with your signs

OpenStudy (anonymous):

ugh. I cant seem to understand it. im pulling my hair out. cause I did that part but then I just get lost.

OpenStudy (anonymous):

@jackellyn

OpenStudy (jackellyn):

Just go step by step 9= 12 +5x -(6x)-7 9-12+7=5x-6x 4=-x x=-4

OpenStudy (anonymous):

"9-12+7=5x-6x" why is it 9-12+7? i thought it would be 9+12-7

OpenStudy (anonymous):

@jackellyn

OpenStudy (jackellyn):

Well the -7 is on the right side and you need to move it to the left so you need to add 7 to both sides

OpenStudy (anonymous):

ohhhhhhh. i get it. and since 12 is postive you are subtracting 12 from both sides?

OpenStudy (anonymous):

@jackellyn

OpenStudy (jackellyn):

Exactly